OpenStudy (anonymous):

Solving system of equations by elimination -2x-9y=-25 and -4x-9y=-23

OpenStudy (noorik.):

just try to subtract the second equation from the first equation (-2x-9y)-(-4x-9y) = -25 - (-23)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} -2x-9y=-25 &{\color{red}{ \times -2}}\implies &\quad \cancel{4x}+18y=+50\\ -4x-9y=-23&& \cancel{-4x}-9y=-23\\ \hline\\ &&\qquad\quad \square y=\square \end{array}\)

OpenStudy (j2lie):

times negative 2 to the first equation and it will equal 4x+18y=50. Then you cross out the x's and you solve for y.